OpenStudy (anonymous):

Demetrius is a salesman whose base pay plus commissions amounted to $51,162 last year. If his base pay was $26,610 and he made $79,200 in sales last year, what was his rate of commission? A. 34% B. 31% C. 52% D. 65%

OpenStudy (anonymous):

commission +base pay =51,162 commission=51,162-26,610 commission=24552 rate= (24552/79200) x 100 solve